OpenStudy (anonymous):

a retailer is selling all 35$ area rugs at 25% off ann is buying a rug and has a coupon for an additional 10% off the sale price. what will she pay for rug?

OpenStudy (anonymous):

i got 12.25

OpenStudy (anonymous):

plz am i right?

OpenStudy (anonymous):

(35*0.75)*0.90 = $23 .625

OpenStudy (anonymous):

dont understand

OpenStudy (anonymous):

oh u have to make the percent into decimals first

OpenStudy (anonymous):

If an item costing $100 is sold for 25% off, then the price is $100-$25 = $75. $100*75% = 100*0.75 = 75.

OpenStudy (turingtest):

yes, first take the original price of the rug to be x the discount (before the coupon) is then 25% of x=0.25x so the price Ann will pay (still before the coupon) is \[x-0.25x=0.75x\]good so far?

OpenStudy (anonymous):

yes i understand now

OpenStudy (anonymous):

to learn math i have to have detailed explanations sometimes thx to u both =)

OpenStudy (turingtest):

so you understand how to get from that price to the price after using the coupon?

OpenStudy (turingtest):

I like to think methodically as well. Math is cumulative and needs to be done step-by-step. Over-complicating things never helped anyone.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats is right thx again for all the help

OpenStudy (turingtest):

Einstein once said "everything should be made as simple as possible, but not simpler." I love that quote.

OpenStudy (anonymous):

hey i love that quote as well an its soooo true!